# Liquid: Language Models are Scalable Multi-modal Generators

### v2

@@ 3.2.1 Comparison with Mainstream LLMs
Method BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c MMLU
Mistral 7B 84.7 83.0 - 81.3 75.3 80.0 55.5 60.1
8$\times$7B - 83.6 - 84.4 77.2 83.1 59.7 70.6
LLAMA2 7B 77.4 78.8 48.3 77.2 69.2 75.2 45.9 45.3
13B 81.7 80.5 50.3 80.7 72.8 77.3 49.4 54.8
34B 83.7 81.9 50.9 83.3 76.7 79.4 54.5 62.6
Chameleon 7B 81.4 79.6 57.0 74.2 70.4 76.1 46.5 52.1
34B 86.0 83.3 63.3 82.7 78.5 84.1 59.7 65.8
Liquid-7B 7B 81.0 81.0 46.7 76.1 72.7 75.6 49.0 56.0
3.2.2 Scaling Results on Language Tasks
To investigate whether visual generation capabilities impact linguistic abilities, we compare the performance of models trained with 30M language data and those trained with a mix of 60M multimodal data on language tasks across different sizes. For language ability, we report the average average scores on 5 language tasks, including HellaSwag, WinoGrande, ARC-Easy, ARC-Challenge, and BoolQ. As shown in Fig. 5, we find a trade-off phenomenon when both tasks were mixed in the training of smaller models; however, as the model size increased, this trade-off gradually disappeared. This confirms that larger models possess sufficient capacity to handle both visual and language spaces generation concurrently.
3.3 Does Language Ability Limit Visual Generation Performance?
In language tasks, we observe a trade-off in performance of multimodal LLMs, raising the question of whether this trade-off also exists in visual generation tasks. To address this question, we compare LLMs of different sizes that have been trained exclusively with text-to-image data and with a mix of multimodal data. Both training utilize the same visual generation data, with the only difference being whether the models retain their linguistic capabilities.
As illustrated in Fig. 6, mixed training leads to higher validation loss on visual generation tasks for models of every size. However, the detrimental effect on the VQA Score decreases as the model size increases. This reduction is partly due to the increased capacity of larger models and may also be attributed to the evaluation metrics approaching saturation. Nevertheless, it can be argued that as the size of the model increases, the trade-off between visual generation and language tasks gradually diminishes or even disappears. This encourages leveraging the scaling capabilities of LLMs into multimodal models.
3.4 How Does Liquid Perform on Visual Understanding Tasks?
After the pre-training stage, the model has already attained substantial image generation capabilities and a decent level of image understanding. To evaluate the visual understanding capabilities, we use 1M LMSYS [75] as text-only instruction data, coupled with 1M text-to-Image data sampled from high-quality data, and 1.5M multi-modal instruction tuning data introduced in Minigemini [33]. This compiles a 3.5M hybrid instruction tuning data for further refining our model. We report results on widely-adopted zero-shot image-based benchmarks, which include VQA-v2 [20], GQA [23], TextVQA [55], POPE [32], and MME [16].
Table 4: Comparison with leading methods on visual language benchmarks. * indicates that images in the training split of these datasets are observed during training. “Und.” and “Gen.” denote “understanding” and “generation”. Our performance surpasses most models that unify understanding and generation, and it is comparable with models dedicated to visual understanding in some tasks. $\dagger$ has a longer pre-training phase, can further enhance its performance.
@@ 3.4 How Does Liquid Perform on Visual Understanding Tasks?
As shown in Tab. 4, compare with the MLLMs with discrete visual token, Liquid outperforms models with stander VQVAE like LWM [40], Chameleon [60], and Show-o [71], and achieves comparable results with VILA-U [70], which have trained a CLIP based multi-codebook VQVAE to improve understanding. However, the performance of MLLMs using discrete visual tokens on visual understanding tasks tends to be lower than mainstream models that employ continuous visual tokens. Most MLLMs with continuous visual token use CLIP features as visual input, it is a significant advantage considering that CLIP is pre-trained on a large-scale image-text pair dataset, leading to a strong alignment between its visual space features and language space features.This alignment substantially aids the LLMs, making it easier for them to understand visual content. In contrast, using image tokens derived directly from VQVAE tokenizer as input means that the corresponding embedding features in the LLM are reinitialized with out any alignment. Without extensive pre-training to align feature spaces, the visual understanding capabilities might be slightly inferior to models using CLIP as visual input. The difference in performance mainly stems from the fact that most VQVAE currently do not align image-text spaces. VILA-U [70] has confirmed that by adding CLIP loss during the VQVAE training to align its visual space, the performance of visual understanding tasks can be improved.
@@
However, we attempt to increase the amount of image-text pairs the model can see during the pre-training phase by training one more epoch on our multi-modal pre-training data, and then we were surprised to discover that the model could achieve better performance on visual understanding tasks. This result indicates that mixed-modality pre-training plays a role similar to CLIP pre-training, aligning text and visual embedding spaces. More pret-raining or more suitable embedding initialization methods could further boost the performance of discrete visual tokens on visual understanding tasks.
As observed in above experiments on visual understanding tasks, extended pretraining enhances comprehension abilities. Notably, during the pretraining process, 80% of the data was trained in a text-to-image format. This observation naturally leads us to question whether increased training with text-to-image data could improve performance on understanding tasks. Furthermore, we explore whether there is a synergistic relationship between these two tasks.
We conduct three sets of experiments. In the first set, we used a combination of 10M text-only data, 10M visual generation data, and 10M visual understanding data, resulting in a total of 30M data for the pre-training phase. The visual understanding data is derived by reversing the visual generation data for the captioning task, and the loss is calculated only on the text output for visual understanding task. All data are subsets of the data introduced in Section 3. The experiment, with a $1:1:1$ data ratio of the three tasks, serves as the baseline. Building on this baseline, we separately add an additional 10M visual generation data and 10M visual understanding data, forming a total of 40M data with data ratios of $1:2:1$ and $1:1:2$, respectively, labeled as ‘Add T2I’ and ‘Add I2T’. We train three GEMMA-7B models using default parameters and evaluate their performance on visual understanding and generation tasks to observe the impact of different data additions on the performance of all tasks. For the visual understanding tasks, after the pre-training phase, we perform the same instruction tuning training and then test them on standard visual understanding benchmarks.
Table 5: The impact between visual understanding and generation tasks. “Visual Gen." refers to the data used for training text-guided image generation, while“Visual Und." refers to the data used for training visual understanding capabilities, specifically in the form of captioning. Compared to the baseline, adding more visual understanding data enhances the visual generation capability, improving the semantic consistency between the generated content and the prompt. Conversely, increasing the visual generation data similarly aids in enhancing the model’s visual understanding ability. This indicates that when the tokens for visual generation and understanding are unified, they share a common optimization objective and can mutually enhance each other.
Method Training Data VIsual Generation Visual Understanding
Text-only Visual Gen. Visual Und. Basic Advanced Overall VQAv2 GQA TextVQA POPE MME
Baseline 10M 10M 10M 0.63 0.58 0.60 60.7 51.3 39.2 75.9 909.6
Add T2I 10M 20M 10M 0.78 0.63 0.69 64.5 53.2 39.8 78.0 1066.8
Add I2T 10M 10M 20M 0.73 0.62 0.66 63.5 53.7 40.5 76.8 1035.1
As shown in Tab. 5, we observe that adding more visual understanding data during pre-training significantly improves the performance of visual generation tasks. Conversely, adding more visual generation data further enhances the performance of visual understanding tasks. This important phenomenon indicates that when the modality spaces for visual understanding and generation are unified, the training of these two tasks can mutually benefit each other. Intuitively, when images are represented using embeddings from the same space, both visual generation and visual understanding tasks require stronger consistency constraints and sufficient interaction between language and visual information. Therefore, the optimization directions for these two tasks are very similar. This further demonstrates the potential of LLMs as general-purpose multi-modal generators. The continuous addition of visual training data can simultaneously enhance the models’ capabilities in both multimodal understanding and generation.
@@ 4 Related Work
Unified Multimodal Understanding and Generation Several early efforts have explored how to construct a unified multi-modal large model for visual generation and understanding based on LLMs. The central challenge lies in tokenizing images into sequence inputs for the LLMs and detokenizing the sequential output of the LLMs back into images, the choice of image tokenizer. Some methods [18, 19, 59, 58] use vision encoders based on ViT like CLIP to encode images into continuous feature maps. The continuous visual space from CLIP can retain more visual information and have a pre-trained, aligned space with language feature. However, the continuous feature often necessitates an additional diffusion module for image detokenization. Other works  [40, 60, 70] employ VQVAE to encode images into discrete tokens and train LLMs to predict them. Still other works [41, 71, 69] use both ViT and VQVAE as tokenizers to garner their benefits. Discrete image features can share the same embedding space with text input, permitting joint reasoning over both modalities within a unified architecture without the requirement for modality-specific components. It is beneficial for model scale-up. Consequently, in our work, we choose VQVAE as the sole image tokenizer. Our work is most similar to LWM [40] and Chameleon [60]. However, they display inferior image understanding and generation capabilities, and need extensive large scale multi-modal pre-training, which is a significant burden. In contrast, we propose to start from any existing LLMs and enhance their visual understanding and generation abilities by continuing training with a small amount of high-quality data, without altering any model structures.

### v4

@@ 4.1 Quantitative Results on Visual Generation
4.2 Comparison with Mainstream LLMs
To validate whether acquiring image understanding and generation capabilities has any impact on the original language abilities of the LLMs, we compare our mixed multimodal pretrained model against other state-of-the-art large language models and multi-modal language models across a suite of popular benchmarks that measure commonsense reasoning and reading comprehension capabilities: HellaSwag [80], WinoGrande [56] , ARC-Easy [10], ARC-Challenge [10], OpenBookQA [44], PIQA [3], SIQA [57], and BoolQ [9]. We also perform an evaluation of the 5-shot results on MMLU [20], a comprehensive benchmark that measures world/in-domain knowledge and problem-solving skills across 57 subjects.
As shown in Tab. 5, Liquid outperforms the well-established language model Llama2 [70] and the mix-pretrained multi-modal language model Chameleon [64] in most tasks, exhibiting undegraded linguistic capabilities. Compared with Chameleon [64], which is mixed pretrained with an extremely large scale of data, Liquid performs training from existing LLMs that already possess decent language capabilities, maintaining these capabilities without degradation. This result validates the efficiency of our training framework and demonstrates that with this framework, we can extend the visual generation and understanding capabilities to LLMs of any structure and size.
@@ 4.2 Comparison with Mainstream LLMs
[…]
Table 6: Comparison with leading methods on visual language benchmarks. * indicates that images in the training split of these datasets are observed during training. “Und.” and “Gen.” denote “understanding” and “generation”. Our performance surpasses most models that unify understanding and generation, and it is comparable with models dedicated to visual understanding in some tasks. $\dagger$ has a longer pre-training phase, $\ddagger$ indicates the use of a VQVAE aligned with CLIP semantics during training.
4.3 Quantitative Results on Visual Understanding
To evaluate the visual understanding capabilities, we use 1M LMSYS [81] as text-only instruction data, coupled with 1M text-to-Image data sampled from high-quality data, and 1.5M multi-modal instruction tuning data introduced in Minigemini [33]. This compiles a 3.5M hybrid instruction tuning data for further refining our pretrained model. We report results on widely-adopted zero-shot image-based benchmarks, which include VQA-v2 [19], GQA [22], TextVQA [59], POPE [32], and MME [16].
@@ 4.3 Quantitative Results on Visual Understanding
As shown in Tab. 6, compare with the MLLMs with discrete visual token, Liquid outperforms models with stander VQVAE like LWM [40], Chameleon [64], and Show-o [76]. However, the performance of MLLMs using discrete visual tokens on visual understanding tasks tends to be lower than mainstream models that employ continuous visual tokens. Most MLLMs with continuous visual token use CLIP features as visual input, it is a significant advantage considering that CLIP is pre-trained on a large-scale image-text pair dataset, leading to a strong alignment between its visual space features and language space features. This alignment substantially aids the LLMs, making it easier for them to understand visual content. In contrast, using image tokens derived directly from VQVAE tokenizer as input means that the corresponding embedding features in the LLM are reinitialized with out any alignment. Without extensive pre-training to align feature spaces, the visual understanding capabilities might be slightly inferior to models using CLIP as visual input. The difference in performance mainly stems from the fact that most VQVAE currently do not align image-text spaces. VILA-U [75] has confirmed that by adding CLIP loss during the VQVAE training to align its visual space, the performance of visual understanding tasks can be improved.
@@
In another hand, we attempt to increase the training step the pre-training phase by add one more epoch, and then we were surprised to discover that the model could achieve better performance on visual understanding tasks (marked with $\dagger$). This result indicates that mixed-modality pre-training plays a role similar to CLIP pre-training, aligning text and visual embedding spaces. More pret-raining or more suitable embedding initialization methods could further boost the performance of discrete visual tokens on visual understanding tasks. To further explore the potential of this paradigm, we replace the VQVAE in Chameleon with UniTok [43] that aligns visual and language spaces during training (marked with $\ddagger$), outperforming VILA-U [75], which trained a CLIP-based multi-codebook VQVAE to improve understanding, and achieving results on par with LLaVA. This demonstrates the importance of visual-semantic space alignment for understanding tasks. Larger-scale pretraining or introducing semantic-aligned priors for visual tokens proves crucial for enhancing the comprehension capabilities of unified multimodal models, representing a key direction for future tokenizer improvements.
4.4 In-Context Learning Across Modalities
Figure 9: Liquid has good in-context learning capability. We feed two image-text pairs and a third image as the context to prompt the model.
In-context learning (ICL) is a hallmark capability of LLMs that enables few-shot adaptation through task demonstrations — has revolutionized how models generalize to unseen scenarios without parameter updates. To investigate whether our unified multimodal architecture Liquid can extend this emergent capability to the vision-language space, we attempt to provide the model with several cross-modal task examples. We find that incorporating interleaved multimodal data [84] into the pretraining dataset enables the model to further learn interactions between multimodal contents, as shown in Fig. 9, given “<image1> is sunny, <image2> is rainy, <image3>”, Liquid correctly associates the third image with “is snowy” by recognizing the weather-related visual-textual pattern. Similarly, when shown images of a cheetah, eagle, and dolphin paired with their motion domains “fast on land”, “fast in air”, the model infers “fast in water” by aligning visual semantics with spatial attributes. These results demonstrate that the model not only grounds individual modalities but also discovers structured cross-modal relationships from limited context, mimicking the compositional reasoning of pure-text ICL. Crucially, this capability emerges without explicit multi-modal alignment supervision, suggesting that the model internalizes a unified representation space where visual and textual patterns cohere into reusable "prompts". Our findings position multi-modal ICL as a promising direction for few-shot adaptation in vision-language systems.
4.5 Visual Comparative Analysis
Impact of Classifier-free Guidance. Classifier-Free Guidance (CFG) scale is hyperparameter that control the trade-off between sample quality and diversity in conditional generative models. The visual variations of generated images with different CFG scales t are illustrated in Fig. 10 As observed, higher CFG scales lead to better alignment between the generated images and the text prompts, but cause more chaotic object structures, stronger stylization, and worse photorealism. For example, when CFG=15, the structure of the book in the image becomes disordered. Conversely, lower CFG scales result in poorer consistency between the image content and the prompt, but improve the photorealism and fine-grained texture details.
Figure 10: The prompt for the first row of images is "A book with glowing runes floating beside a mystic crystal." The prompt for the second row is "A bird nocks an arrow." Higher CFG scales enhance the consistency between the model and the text prompt but degrade photorealism, while lower CFG scales improve image realism at the cost of weaker semantic alignment.
Figure 11: Visual generation comparison between Liquid and other unified multim-odal models for understanding and generation
Comparison with Other Models. In Fig. 11, we present a comparison between Liquid and other unified multi-modal large models in terms of visual generation quality. Compared to models based on discrete multi-codebook (VILA-U), diffusion processes (Show-o), and multimodal tokenizers (Janus), Liquid demonstrates superior performance in knowledge-aware image generation (first row), scene generation accuracy and small-scale facial details (second row), as well as structural coherence of objects (third row, vehicles). This visual comparison quantitatively validates Liquid’s advancements in generating high-fidelity, semantically consistent images while maintaining strong multi-modal understanding capabilities.
4.6 Discussion
Liquid addresses the following limitations of previous works:
1. Previous unified multimodal models usually suffered from degraded language capabilities, limiting their broader applicability. Liquid demonstrates that a unified multimodal model can maintain on-par language performance even after continued training, preserving its potential as a versatile foundation model.
2. No prior work has explored whether LLMs retain the power-law scaling laws observed in language tasks when extended to visual generation tasks. We prove this alignment and further show that vision can be effectively learned by LLMs as a form of language.
3. Previous works [47, 74] observed conflicts between visual understanding and generation tasks. We discover that the unified token space enables visual generation and comprehension tasks to mutually enhance each other, effectively removing the conflict.
@@ 5 Related Work
Unified Multimodal Understanding and Generation Several early efforts have explored how to construct a unified multi-modal large model for visual generation and understanding based on LLMs. The central challenge lies in tokenizing images into sequence inputs for the LLMs and detokenizing the sequential output of the LLMs back into images, the choice of image tokenizer. Some methods [17, 18, 63, 62] use vision encoders based on ViT like CLIP to encode images into continuous feature maps. The continuous visual space from CLIP can retain more visual information and have a pre-trained, aligned space with language feature. However, the continuous feature often necessitates an additional diffusion module for image detokenization. Other works  [40, 64, 75] employ VQVAE to encode images into discrete tokens and train LLMs to predict them. Still other works [41, 76, 74, 49] use both ViT and VQVAE as tokenizers to garner their benefits. Discrete image features can share the same embedding space with text input, permitting joint reasoning over both modalities within a unified architecture without the requirement for modality-specific components. It is beneficial for model scale-up. Consequently, in our work, we choose VQVAE as the sole image tokenizer. Our work is most similar to LWM [40] and Chameleon [64]. However, they display inferior image understanding and generation capabilities, and need extensive large scale multi-modal pre-training, which is a significant burden. In contrast, we propose to start from any existing LLMs and enhance their visual understanding and generation abilities by continuing training with a small amount of high-quality data, without altering any model structures.